OpenStudy (anonymous):

Solve the following equation, answer as a reduced, mixed number.3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1)

Nnesha (nnesha):

first you have to distribute bracket can you do that ??

Nnesha (nnesha):

do you know how to distribute ???

OpenStudy (dtan5457):

Yes. Distribute out all brackets. You get 3x-6+8x-24=6x-24+16x+8 add up like terms to get 11x-30=22x+16 get x alone -11x=46 divide by -11 x=-4 2/11

Nnesha (nnesha):

@dtan5457 is that really hlpful ??

OpenStudy (anonymous):

wrong dtan

Nnesha (nnesha):

hope you know that direct answer not allowed

OpenStudy (anonymous):

uh oh.....

OpenStudy (anonymous):

u'ssss in troublleeeee

OpenStudy (anonymous):

but seriously someone help me

OpenStudy (dtan5457):

yeah I see what i did wrong. it's suppose to be 11x-30=22x-16 not positive 16

OpenStudy (dtan5457):

therefore x= -1 3/11

OpenStudy (anonymous):

Well as @Nnesha said, distribute first, then put the like terms together, then solve for the variable on one side. I suggest trying to get thee variable positive to reduce the amount of work.

OpenStudy (anonymous):

@Nnesha hey how u earn engagement points

Nnesha (nnesha):

idk i guess by asking quesiton

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

just cheat for pity sake!

OpenStudy (texaschic101):

thats what is happening here :(

OpenStudy (anonymous):

k then someone cheat and tell me the answer

OpenStudy (dtan5457):

i already did.

OpenStudy (anonymous):

u were wrong but i got it so nvm

OpenStudy (anonymous):

actually the second time u were right sorry